OpenStudy (anonymous):

Find the measure of the smallest angle in triangle abc. Round your answer to the nearest tenth. The legs are 7 and 11 neither of them is a hypotenuse. plz help my calculator doesn't want to work for me.

OpenStudy (anonymous):

Sorry about that. Forgot the drawing.

OpenStudy (anonymous):

that's ok but when i did it that way i came up with 36.479 and so on

OpenStudy (anonymous):

thank here's your medal

OpenStudy (anonymous):

\[N\left[\tan ^{-1}\left(\frac{7}{11}\right),20\right]=0.566729217523506 \]